# Rotationally invariant dynamical lattice regulators for Euclidean quantum field theories

### v1

@@ Abstract
We introduce a dynamical–lattice regulator (DLR) for Euclidean quantum field theories on a fixed hypercubic graph $\Lambda\simeq\mathbb{Z}^{d}$, in which the embedding $x:\Lambda\to\mathbb{R}^{d}$ is promoted to a dynamical field and integrated over subject to shape–regularity constraints. The total action is local on $\Lambda$, gauge invariant, and depends on $x$ only through Euclidean invariants built from edge vectors (local metrics, volumes, etc.), hence the partition function is exactly covariant under the global Euclidean group $\mathrm{SE}(d)$ at any lattice spacing. The intended symmetry–restoring mechanism is not rigid global zero modes but short–range local twisting of the embedding that mixes local orientations; accordingly, our universality discussion is conditioned on a short–range geometry hypothesis (SR): after quotienting the global $\mathrm{SE}(d)$ modes, connected correlators of local geometric observables have correlation length $O(1)$ in lattice units.
We prove Osterwalder–Schrader reflection positivity for the coupled system with embedding $x$ and generic gauge/matter fields $(U,\Phi)$ in finite volume by treating $x$ as an additional multiplet of scalar fields on $\Lambda$. Assuming (SR), integrating out $x$ at fixed cutoff yields a local Symanzik effective action in which geometry fluctuations generate only $\mathrm{SO}(d)$–invariant irrelevant operators (and finite renormalizations); in particular, in $d=4$ we recover the standard one–loop $\beta$–function in a scalar $\phi^{4}$ test theory. Finally, we describe a practical local Monte Carlo update and report $d=2$ proof–of–concept simulations showing a well–behaved geometry sector and a substantial reduction of axis–vs–diagonal cutoff artefacts relative to a fixed lattice at matched bare parameters.
@@ 1.1 Motivation
The aim of this paper is to develop and analyse a dynamical lattice regulator (DLR) for Euclidean QFT that combines the robust structural features of standard lattice gauge theory (locality, exact gauge invariance, Osterwalder–Schrader reflection positivity, and a fixed combinatorial graph) with an exactly Euclidean–symmetric treatment of geometry already at finite lattice spacing. Concretely, we fix once and for all an abstract hypercubic lattice
$\Lambda:=\mathbb{Z}^{d}$ (1.1)
(or its finite periodic version for numerical work), which carries the usual gauge and matter fields. We then promote the embedding of lattice sites into $\mathbb{R}^{d}$ to a dynamical field
@@
and it is compatible with the specific reflection involution used in the OS positivity proof. However, global covariance alone is too weak to address the usual lattice concern: it would also hold if one only sampled rigidly rotated/translated copies of a regular embedding. The distinctive feature of the present regulator is instead local twisting: the local frame orientation varies from site to site with short correlation length, producing local orientation averaging and thereby suppressing direction-dependent cutoff effects. In effect, the embedding field induces fluctuating local couplings for the gauge and matter fields while never changing the underlying hypercubic adjacency: neighbours in $\Lambda$ remain neighbours, even as their physical separation fluctuates.
@@
  • is expected to lie in the same universality class as standard hypercubic lattice regulators: perturbatively we verify this at one loop, and in $d=2$ we find matching critical scaling between baseline and dynamical ensembles (Section 5.6); more generally we formulate a universality statement as Conjecture 4.8 under the short-range geometry hypothesis (SR).
The random lattice literature has shown that geometric disorder can soften or eliminate roughening transitions. While modern Monte Carlo methods operate successfully within the rough phase, we anticipate that the dynamical-lattice regulator will similarly mitigate the lattice anisotropy that “pins” topological objects to the grid axes; verifying the softening of this transition would serve as a powerful diagnostic of rotational symmetry restoration. Furthermore, while the fixed topological graph implies that the fermion doubling problem will likely persist ([23]), averaging over embeddings may significantly reduce direction-dependent artefacts (such as taste splitting) in fermionic correlators at finite cutoff, in a way that is closer in spirit to random-lattice approaches but technically much easier to interface with existing hypercubic lattice codes. A systematic study of these phenomena is left to future work.
@@ 1.2 Overview of the construction
We work with an abstract, fixed hypercubic lattice $\Lambda$ (typically a discrete torus in numerical applications). The usual gauge and matter fields live on $\Lambda$ exactly as in Wilson’s formulation: link variables $U_{\mu}(n)\in G$ on oriented nearest–neighbour links, and matter fields $\Phi(n)$ in some representation of $G$ on the sites. The new ingredient is a geometry field
@@
which specifies the physical position $x(n)\in\mathbb{R}^{d}$ of each lattice site. From $x$ we construct local edge vectors, metrics and cell volumes, and we restrict $x$ to an admissible subset $\mathcal{X}_{\mathrm{adm}}$ of $(\mathbb{R}^{d})^{\Lambda}$ by imposing standard shape–regularity conditions. The admissible set is chosen so that all local geometries are uniformly comparable to a regular hypercubic mesh.
@@
where $S_{x}[x]$ is a geometry action and $S_{\mathrm{fields}}$ is the usual gauge–matter action written in terms of the local metric and volume induced by $x$. Both pieces are built from Euclidean invariants of the embedding data and from gauge–invariant combinations of the fields. In particular, the resulting theory on $\mathbb{R}^{d}$ is globally covariant under global translations and proper rotations, that is, under the special Euclidean group $\mathrm{SE}(d)=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$; moreover, it is compatible with the discrete reflection involution used in the Osterwalder–Schrader positivity argument, cf. [15, 16]. This statement concerns the unfixed formulation; in practice one may fix the global translation/rotation zero modes without affecting covariant expectations or the local twisting diagnostics.
Standard Wilson lattice gauge theory is recovered as a stiff–geometry limit of the dynamical lattice regulator in which $S_{x}[x]=+\infty$ unless $x=x_{\mathrm{reg}}$, where $x_{\mathrm{reg}}(n)=a\,n$ is a regular embedding. This limit is not Euclidean–invariant: freezing the embedding selects a preferred orientation and breaks the continuous Euclidean symmetry explicitly down to the hypercubic symmetry group. In particular, merely averaging over rigid global rotations of this frozen embedding does not change the local stencil anisotropy; improving isotropy requires fluctuations that mix orientations locally. The novelty is that the geometry remains annealed in a way that preserves locality and the transfer-matrix/OS-positivity framework on the fixed hypercubic graph, while enabling local orientation mixing that can suppress hypercubic anisotropies; the stiff (Wilson) limit is recovered by sending the geometry stiffness to infinity.
For conceptual questions (Euclidean covariance, reflection positivity, continuum limit) it is most convenient to formulate the theory on $\mathbb{R}^{d}$, with $\Lambda$ thought of as an infinite hypercubic lattice and $x(n)\in\mathbb{R}^{d}$. For finite–volume simulations we instead interpret $\Lambda$ as a discrete torus and take $x(n)\in\mathbb{T}^{d}$, using shortest–distance representatives to define edge vectors; this torus version is described in §2.4. In both cases the underlying combinatorial graph is kept fixed, while the embedding $x$ fluctuates subject to admissibility constraints.
@@ 1.3 Main results and scope of this paper
We now summarize the main theoretical and practical conclusions.
@@
  • Global Euclidean covariance and reflection positivity. We show that if the geometry and field actions are built from local Euclidean invariants of the edge vectors and from gauge–invariant field combinations, then the dynamical–lattice theory on $\mathbb{R}^{d}$ is exactly invariant under the special Euclidean group $\mathrm{SE}(d)$ at any fixed lattice spacing. We distinguish this formal global covariance from the physically relevant notion of reduced direction-dependent cutoff artefacts, which we probe numerically and relate to short-range local twisting. Moreover, by viewing the geometry field $x$ simply as an additional set of scalar fields on the abstract lattice, we prove Osterwalder–Schrader reflection positivity for a broad class of gauge and matter actions. The transfer–matrix and Hilbert–space reconstruction therefore go through in close analogy with standard lattice gauge theory [21, 22].
@@
  • Locality, admissibility and the geometry sector. We define an admissible set $\mathcal{X}_{\mathrm{adm}}$ of embeddings by local bounds on edge lengths, angles and orientation, in such a way that the local metrics and volumes are uniformly comparable to those of a regular hypercubic lattice. We discuss the global structure of $\mathcal{X}_{\mathrm{adm}}$, the choice of principal component, and simple local Monte–Carlo moves for updating $x$ that preserve admissibility while remaining ergodic within this component.
@@
  • Perturbative universality. As a first perturbative check we analyze a scalar $\phi^{4}$ theory on a dynamical lattice. In the Symanzik framework [15, 16], and assuming the geometry sector is in a short–range (massive) regime, integrating out the embedding fluctuations yields a local $\mathrm{SE}(d)$-invariant effective action: the relevant and marginal operators are therefore the same as in the continuum, while cutoff effects appear only through higher–dimensional $\mathrm{SE}(d)$-scalar operators. In particular, $\mathrm{H}(d)$-invariant but non-$\mathrm{SO}(d)$-invariant counterterms that are permitted on a fixed hypercubic lattice are forbidden here by symmetry. Expanding around a regular embedding as $x(n)=an+\xi(n)$ and treating the geometry fluctuations $\xi$ as auxiliary fields, we show that the induced interactions are local and, conditional on the short-range geometry hypothesis (SR), irrelevant in the renormalization-group sense. In particular, at one loop the running of the quartic coupling agrees with that of the standard hypercubic discretization. This is naturally phrased in the Symanzik effective-action language: the geometry sector can renormalize the bare-to-renormalized map and generate only $\mathrm{SO}(d)$-scalar higher-dimension operators, but does not introduce new relevant/marginal directions beyond those allowed by the target continuum symmetries; see Conjecture 4.8.
@@
  • Monte–Carlo implementation and numerical tests. We describe a first proof of concept Monte–Carlo implementation in $d=2$ for a scalar theory on a periodic lattice, using local Metropolis updates for both fields and geometry. Basic observables (masses, susceptibilities) agree with those from a static lattice at the same bare parameters within statistical errors, and simple diagnostics of the geometry sector (edge–length, angle and volume distributions) show no pathological concentration on highly distorted meshes. These tests support the claim that the regulator is numerically viable and lies in the expected universality class.
@@
The rest of the paper is organised as follows. In Section 2 we develop the geometry sector: admissible embeddings, local shape-regularity, and Monte–Carlo moves. Section 3 introduces the field actions, demonstrates exact $\mathrm{E}(d)$ invariance, and presents explicit scalar, gauge, and fermion examples. Section 4 establishes reflection positivity and discusses the Symanzik effective theory and continuum universality. Section 5 describes our initial numerical experiments, and Section 6 summarises our findings and outlines possible extensions.
@@ 1.4 Comparison to existing approaches
In the random lattice (RL) approach, one typically samples a random set of points in a domain $\Omega\subset\mathbb{R}^{d}$ (often from a Poisson process), constructs a Voronoi tessellation and its dual Delaunay triangulation, and discretizes the field theory on the resulting random cell complex, cf. [4, 5]. As the points move, the connectivity of the lattice (the adjacency graph) changes discontinuously when Voronoi cells flip; the adjacency matrix is a discontinuous function of the site positions. In this setting the geometry is random, but its law is largely specified by an external stochastic prescription for the points, and reflection positivity is hard to control.
The “pseudorandom lattice” construction of Colangelo et al. [6] also aims at improving rotational invariance, but does so while preserving the hypercubic graph: one draws exactly one site inside each elementary hypercubic cell of a reference lattice and connects sites belonging to nearest–neighbour cells. Thus the combinatorial structure and sparsity pattern are the same as on a standard hypercubic lattice, while the geometric embedding (link lengths and angles) is quenched random data chosen once from a tuned probability distribution. This improves the statistical isotropy of simple geometric observables, but does not enforce shape regularity or exact $\mathrm{SE}(d)$ symmetry at finite lattice spacing, and the random geometry does not enter as a dynamical field in the same path integral as the gauge and matter sectors.
By contrast, in the dynamical–lattice regulator the abstract hypercubic topology is fixed once and for all. The adjacency relation on $\Lambda$ does not depend on the embedding $x$, and moving the sites never creates or destroys links. What fluctuates is only the embedding
@@
which is treated as a genuine dynamical field with a local action $S_{x}[x]$ in the same path integral as the gauge and matter fields. Geometric fluctuations are therefore “annealed” (dynamical) rather than quenched (frozen): the randomness of the geometry is intrinsic to the quantum Lagrangian. (Note that we use “annealed” here to refer to the spacetime lattice itself, distinct from the “quenched approximation” regarding fermion loops).
@@
Regge calculus keeps a fixed simplicial connectivity and promotes the edge lengths to dynamical variables, with an action that discretizes the Einstein–Hilbert functional. In dynamical triangulations (DT) and causal dynamical triangulations, one takes a complementary viewpoint: the edge lengths are kept fixed (typically to a few discrete values) and one instead sums over triangulations themselves, so that the gluing pattern of simplices becomes the dynamical variable, again weighted by a Regge–type action plus matter contributions. In both cases the geometric degrees of freedom are fundamental and the primary goal is to describe quantum gravity [7, 8, 9, 10].
Structurally, the dynamical–lattice regulator is closest to Regge calculus in that it uses a fixed abstract complex together with fluctuating geometric data. There are, however, two important differences. First, the basic variables here are the embedding coordinates $x(n)$, from which edge lengths and local metrics are derived; the geometry sector is introduced as an auxiliary regulator field whose role is to restore $\mathrm{SE}(d)$ symmetry and potentially reduce lattice artefacts, not to encode an independent gravitational degree of freedom. Second, the construction is deliberately tied to a hypercubic abstract lattice so as to maintain the locality structure and transfer–matrix decomposition of standard lattice gauge theory, which is crucial for the reflection–positivity argument and for reusing existing simulation technology. One may view dynamical–lattice as a kind of “Regge–like” dynamics for a fixed hypercubic mesh, with an action tuned for regularization rather than for quantum gravity.
@@
Modern lattice gauge theory extensively uses “smearing” techniques that replace the link variables by local averages, as in APE, Stout, or HYP smearing, cf. [11, 12, 13], and the gradient flow that defines a diffusion in field space [14]. These methods replace the oscillating gauge links with locally averaged variables in the definition of operators or the action. While this shares our goal of mitigating UV noise, the mechanism is fundamentally different: smearing smooths the fields over a fixed, rigid geometry, whereas the dynamical-lattice regulator smooths the geometry itself. By annealing the metric, our approach potentially allows the lattice to accommodate topological objects that might otherwise be frozen by the rigidity of a static grid, a feature that algorithmic smoothing of fields cannot replicate. Equivalently, DLR can be viewed as an annealed geometric averaging (or “geometric smearing”): instead of diffusing/smoothing the dynamical fields on a fixed hypercubic background, we diffuse the coordinate/metric degrees of freedom seen by the fields by integrating over embeddings. A key difference is that the $\mathrm{SE}(d)$ symmetry is exact at the level of the regulator and partition function (modulo the global zero modes), whereas smearing/flow improve operators or effective descriptions at finite flow time but do not enlarge the underlying symmetry group of the static lattice regulator beyond the hypercubic subgroup. In this sense DLR realises intrinsic stochastic rotational averaging inside the path integral, rather than an external rotation/projection applied to measured data.
@@
Standard lattice regulators rely on the Symanzik improvement program to systematically remove discretization errors by adding higher-dimension irrelevant operators to the action [15, 16]. On a static hypercubic lattice, the breaking of the Euclidean group $SE(d)$ down to the hypercubic subgroup $H(d)$ causes single continuum operators to split into multiple lattice operators (e.g., the Laplacian square $(\partial^{2})^{2}$ splits into rotationally invariant and non-invariant parts). Restoring rotational symmetry to high order therefore requires tuning the coefficients of multiple non-invariant counterterms. In contrast, because the dynamical-lattice regulator preserves exact $SE(d)$ invariance at finite cutoff, the effective action is constrained to contain only global scalars. This potentially simplifies the improvement program: although the field content is enlarged by the geometry variables, the basis of dangerous irrelevant operators is much smaller. Non-scalar counterterms are forbidden by exact symmetry, meaning one need only tune the coefficients of $E(d)$-invariant operators to achieve higher-order accuracy.
Moving–mesh methods.
From the perspective of numerical analysis, the construction is reminiscent of Arbitrary Lagrangian–Eulerian (ALE) and other moving–mesh formulations, in which one works with a fixed reference mesh and a time-dependent embedding into physical space, subject to shape–regularity constraints [17, 18]. The crucial difference is that in standard ALE methods the mesh motion is prescribed by a deterministic algorithm (often to equidistribute discretization error), and the mesh is not part of the dynamical system. In the dynamical–lattice regulator the embedding variables $x(n)$ are integrated over quantum–mechanically; one may regard the scheme as a “quantum ALE” version of lattice field theory.
These comparisons place the dynamical–lattice regulator as a conservative middle ground between rigid hypercubic lattices and more radical fluctuating–topology approaches: the topology and constructive–QFT structure of Wilson’s lattice theory are retained, while an intrinsic, dynamically fluctuating geometry is used to restore global Euclidean symmetry and, potentially, to reduce lattice artefacts.
2 Geometry of the dynamical–lattice regulator

### v2

@@ Abstract
We introduce a dynamical lattice regulator for Euclidean quantum field theories on a fixed hypercubic graph $\Lambda\simeq\mathbb{Z}^{d}$, in which the embedding $x:\Lambda\to\mathbb{R}^{d}$ is promoted to a dynamical field and integrated over subject to shape regularity constraints. The total action is local on $\Lambda$, gauge invariant, and depends on $x$ only through Euclidean invariants built from edge vectors (local metrics, volumes, etc.), hence the partition function is exactly covariant under the global special Euclidean group $\mathrm{SE}(d)$ at any lattice spacing. The intended symmetry restoring mechanism is not rigid global zero modes but short–range local twisting of the embedding that mixes local orientations. Our universality discussion is conditioned on a short-range geometry hypothesis (SR): after quotienting the global $\mathrm{SE}(d)$ modes, connected correlators of local geometric observables have correlation length $O(1)$ in lattice units.
We prove Osterwalder–Schrader reflection positivity for the coupled system with embedding $x$ and generic gauge and matter fields $(U,\Phi)$ in finite volume by treating $x$ as an additional multiplet of scalar fields on $\Lambda$. Assuming (SR), integrating out $x$ at fixed cutoff yields a local Symanzik effective action in which geometry fluctuations generate only $\mathrm{SO}(d)$–invariant irrelevant operators and finite renormalizations. For example, in $d=4$ we recover the standard one–loop $\beta$–function in a scalar $\phi^{4}$ test theory. Finally, we describe a practical local Monte Carlo update and report $d=2$ proof–of–concept simulations showing a well behaved geometry sector and a substantial reduction of axis–vs–diagonal cutoff artefacts relative to a fixed lattice at matched bare parameters.
@@ 1.1 Motivation
The aim of this paper is to develop and analyse a dynamical lattice regulator (DLR) for Euclidean QFT that combines the robust structural features of standard lattice gauge theory (locality, exact gauge invariance, Osterwalder–Schrader reflection positivity, and a fixed combinatorial graph) with an exactly (proper) Euclidean–symmetric treatment of geometry already at finite lattice spacing. Concretely, we fix once and for all an abstract hypercubic lattice
$\Lambda=\mathbb{Z}^{d},$ (1.1)
or its finite periodic version for numerical work, which carries the usual gauge and matter fields. We then promote the embedding of lattice sites into $\mathbb{R}^{d}$ to a dynamical field
@@
and it is compatible with the specific reflection involution used in the OS positivity proof. However, the mere fact that the partition function is globally $\mathrm{SE}(d)$–covariant is, by itself, too weak to address the usual lattice concern: the same formal covariance would also hold if the geometry measure were concentrated near globally rotated/translated copies of a single rigid embedding. What matters for suppressing hypercubic cutoff artefacts is how the global symmetry is realised in typical configurations at fixed cutoff.
The operating regime we target is a short-range geometry regime (SR). After fixing or quotienting the global $\mathrm{SE}(d)$ zero modes, typical configurations must exhibit short-range local orientation mixing of the induced frames from cell to cell. In such a regime, global $\mathrm{SE}(d)$ covariance cannot be realised “rigidly”; it must be implemented through local orientation mixing of the induced frames from cell to cell. We refer to this mechanism as local twisting, and it is the one probed in the numerics: neighbours in $\Lambda$ remain neighbours, but their relative physical orientation varies with short correlation length while their physical separation fluctuates within the admissible range.
@@
  • is expected to lie in the same universality class as standard hypercubic lattice regulators: perturbatively we verify this at one loop, and in $d=2$ we find matching critical scaling between baseline and dynamical ensembles; more generally we formulate a universality statement as Conjecture 4.9 under the short-range geometry hypothesis (SR).
From the lattice QCD viewpoint, the dynamical–lattice regulator may be viewed as an annealed geometric averaging in coordinate space, complementary to field–space smoothing techniques such as link smearing and the Wilson flow. In gauge theories we expect such averaging to reduce direction dependent cutoff effects that can “pin” extended or topological structures to the lattice axes, potentially mitigating forms of topological freezing at fixed cutoff. At the same time, because the underlying graph topology remains hypercubic, issues such as the fermion doubling are not expected to be resolved by the regulator; the intended benefit for fermions is instead a reduction of anisotropic artefacts (e.g. taste splitting) through orientation mixing. A systematic study of these gauge theory implications is left to future work.
@@ 1.2 Overview of the construction
We work with an abstract, fixed hypercubic lattice $\Lambda$, which would typically be a discrete torus in numerical applications. The usual gauge and matter fields live on $\Lambda$ exactly as in Wilson’s formulation: link variables $U_{\mu}(n)\in G$ on oriented nearest neighbour links, and matter fields $\Phi(n)$ in some representation of $G$ on the sites. Here $n\in\Lambda$ denotes the sites, and $\mu\in\{1,\ldots,d\}$ the spacial directions. The new ingredient is a geometry field
@@
which specifies the physical position $x(n)\in\mathbb{R}^{d}$ of each lattice site. From $x$ we construct local edge vectors, metrics and cell volumes, and we restrict $x$ to an admissible subset $\mathcal{X}_{\mathrm{adm}}$ of $(\mathbb{R}^{d})^{\Lambda}$ by imposing standard shape regularity conditions. The admissible set is chosen so that all local geometries are uniformly comparable to a regular hypercubic mesh with some scale $a>0$.
@@
where $S_{x}[x]$ is a geometry action and $S_{\mathrm{fields}}$ is the usual gauge–matter action written in terms of the local metric and volume induced by $x$. Both pieces are built from Euclidean invariants of the embedding data and from gauge invariant combinations of the fields. The resulting theory on $\mathbb{R}^{d}$ is globally covariant under translations and proper rotations, i.e., under the special Euclidean group $\mathrm{SE}(d)=\mathrm{SO}(d)\ltimes\mathbb{R}^{d}$. It is also compatible with the discrete reflection involution used in the Osterwalder–Schrader positivity argument, cf. [16, 17]. This statement concerns the unfixed formulation; in practice one may fix the global translation/rotation zero modes without affecting covariant expectations or the local twisting diagnostics.
Standard Wilson lattice gauge theory is recovered as a stiff–geometry limit of the dynamical lattice regulator in which $S_{x}[x]=+\infty$ unless $x=x_{\mathrm{reg}}$, where $x_{\mathrm{reg}}(n)=a\,n$ is a regular embedding. This limit is not Euclidean–invariant: freezing the embedding selects a preferred orientation and breaks the continuous Euclidean symmetry explicitly down to the hypercubic symmetry group. Merely averaging over rigid global rotations of this frozen embedding does not change the local stencil anisotropy; improving isotropy requires fluctuations that mix orientations locally. The novelty is that the geometry remains annealed in a way that preserves locality and the transfer-matrix/OS-positivity framework on the fixed hypercubic graph. This enables local orientation mixing that can suppress hypercubic anisotropies; the stiff (Wilson) limit is recovered by sending the geometry stiffness to infinity.
For conceptual questions (Euclidean covariance, reflection positivity, continuum limit) it is most convenient to formulate the theory on $\mathbb{R}^{d}$, with $\Lambda$ thought of as an infinite hypercubic lattice and $x(n)\in\mathbb{R}^{d}$. For finite volume simulations we instead interpret $\Lambda$ as a discrete torus and take $x(n)\in\mathbb{T}^{d}$, using shortest distance representatives to define edge vectors; this torus version is described in §2.4. In both cases the underlying combinatorial graph is kept fixed, while the embedding $x$ fluctuates subject to admissibility constraints.
@@ 1.3 Main results and scope of this paper
We summarize the main theoretical and practical conclusions obtained in this paper for dynamical lattice regulators (DLR).
@@
  • Global Euclidean covariance and reflection positivity. At fixed lattice spacing the coupled theory is exactly $\mathrm{SE}(d)$–covariant, and we prove Osterwalder–Schrader reflection positivity for a broad class of gauge and matter actions. We distinguish this formal global covariance from the physically relevant notion of reduced direction-dependent cutoff artefacts, which we probe numerically and relate to short-range local twisting.
@@
  • Locality, admissibility and the geometry sector. We define an admissible set $\mathcal{X}_{\mathrm{adm}}$ of embeddings by local bounds on edge lengths, angles and orientation, in such a way that the local metrics and volumes are uniformly comparable to those of a regular hypercubic lattice. We discuss the global structure of $\mathcal{X}_{\mathrm{adm}}$, the choice of principal component, and simple local Monte Carlo moves for updating $x$ that preserve admissibility while remaining ergodic within this component.
@@
  • Perturbative universality. As a first perturbative check we analyze scalar $\phi^{4}$ theory on a dynamical lattice. In the Symanzik framework [16, 17], and assuming the short-range geometry hypothesis (SR), integrating out the embedding fluctuations yields a local $\mathrm{SE}(d)$-invariant effective action: the relevant and marginal operators are therefore the same as in the continuum, while cutoff effects appear only through higher-dimensional $\mathrm{SO}(d)$-scalar operators. Expanding around a regular embedding $x(n)=an+\eta(n)$ and treating $\eta$ as auxiliary fields, we argue that, conditional on (SR), the induced interactions are local and RG-irrelevant. Moreover, at one loop the running of the quartic coupling agrees with that of the standard hypercubic discretization, in line with Conjecture 4.9.
@@
  • Monte Carlo implementation and numerical tests. We describe a first proof of concept Monte Carlo implementation in $d=2$ for a scalar theory on a periodic lattice, using local Metropolis updates for both fields and geometry. Basic observables (masses, susceptibilities) agree with those from a static lattice at the same bare parameters within statistical errors, and simple diagnostics of the geometry sector (edge length, angle and volume distributions) show no pathological concentration on highly distorted meshes. These tests support the claim that the regulator is numerically viable and lies in the expected universality class.
@@
The rest of the paper is organised as follows. In Section 2 we develop the geometry sector: admissible embeddings, local shape-regularity, and Monte Carlo moves. Section 3 introduces the field actions, demonstrates exact $\mathrm{SE}(d)$ invariance, and presents explicit scalar, gauge, and fermion examples. Section 4 establishes reflection positivity and discusses the Symanzik effective theory and continuum universality. Section 5 describes our initial numerical experiments, and Section 6 summarises our findings and outlines possible extensions.
@@ 1.4 Comparison to existing approaches
In the random lattice (RL) approach, one typically samples a random set of points in a domain $\Omega\subset\mathbb{R}^{d}$, often from a Poisson process, constructs a Voronoi tessellation and its dual Delaunay triangulation, and discretizes the field theory on the resulting random cell complex, cf. [4, 5]. As the points move, the connectivity of the lattice changes discontinuously when Voronoi cells flip; the adjacency matrix is a discontinuous function of the site positions. In this setting the geometry is random, but its law is largely specified by an external stochastic prescription for the points, and reflection positivity is hard to control.
The “pseudorandom lattice” construction of Colangelo et al. [6, 7] also aims at improving rotational invariance, but does so while preserving the hypercubic graph: one draws exactly one site inside each elementary hypercubic cell of a reference lattice and connects sites belonging to nearest–neighbour cells. Thus the combinatorial structure and sparsity pattern are the same as on a standard hypercubic lattice, while the geometric embedding (link lengths and angles) is quenched random data chosen once from a tuned probability distribution. This improves the statistical isotropy of simple geometric observables, but does not enforce shape regularity or exact $\mathrm{SE}(d)$ symmetry at finite lattice spacing, and the random geometry does not enter as a dynamical field in the same path integral as the gauge and matter sectors.
By contrast, in DLR the abstract hypercubic topology is fixed once and for all. The adjacency relation on $\Lambda$ does not depend on the embedding $x$, and moving the sites never creates or destroys links. What fluctuates is only the embedding
@@
which is treated as a genuine dynamical field with a local action $S_{x}[x]$ in the same path integral as the gauge and matter fields. Geometric fluctuations are therefore “annealed” (dynamical) rather than quenched (frozen): the randomness of the geometry is intrinsic to the quantum Lagrangian. Note that we use “annealed” here to refer to the spacetime lattice itself, distinct from the “quenched approximation” regarding fermion loops.
@@
Regge calculus keeps a fixed simplicial connectivity and promotes the edge lengths to dynamical variables, with an action that discretizes the Einstein–Hilbert functional. In dynamical triangulations (DT) and causal dynamical triangulations, one takes a complementary viewpoint: the edge lengths are kept fixed (typically to a few discrete values) and one instead sums over triangulations themselves, so that the gluing pattern of simplices becomes the dynamical variable, again weighted by an Einstein–Hilbert–type action plus matter contributions. In both cases the geometric degrees of freedom are fundamental and the primary goal is to describe quantum gravity [8, 9, 10, 11].
Structurally, DLR is closest to Regge calculus in that it uses a fixed abstract complex together with fluctuating geometric data. There are, however, two important differences. First, the basic variables here are the embedding coordinates $x(n)$, from which edge lengths and local metrics are derived; the geometry sector is introduced as an auxiliary regulator field whose role is to restore $\mathrm{SE}(d)$ symmetry and potentially reduce lattice artefacts, not to encode an independent gravitational degree of freedom. Second, the construction is deliberately tied to a hypercubic abstract lattice so as to maintain the locality structure and transfer matrix decomposition of standard lattice gauge theory, which is crucial for the reflection positivity argument and for reusing existing simulation technology. One may view dynamical lattice as a Regge–like dynamics for a fixed hypercubic mesh, with an action tuned for regularization rather than for quantum gravity.
@@
Table 1: Comparison between the random lattice approaches, dynamical triangulations, Regge calculus, and the dynamical lattice regulators.
@@
Modern lattice gauge theory extensively uses “smearing” techniques that replace the link variables by local averages, as in APE, Stout, or HYP smearing, cf. [12, 13, 14], and the gradient flow that defines a diffusion in field space [15]. These methods replace the oscillating gauge links with locally averaged variables in the definition of operators or the action. While this shares our goal of mitigating UV noise, the mechanism is fundamentally different: smearing smooths the fields over a fixed, rigid geometry, whereas DLR smooths the geometry itself. By annealing the metric, our approach potentially allows the lattice to accommodate topological objects that might otherwise be frozen by the rigidity of a static grid, a feature that algorithmic smoothing of fields cannot replicate. Equivalently, DLR can be viewed as an annealed geometric averaging (or “geometric smearing”): instead of diffusing/smoothing the dynamical fields on a fixed hypercubic background, we diffuse the coordinate/metric degrees of freedom seen by the fields by integrating over embeddings. A key difference is that the $\mathrm{SE}(d)$ symmetry is exact at the level of the regulator and partition function (modulo the global zero modes). Smearing/flow improve operators or effective descriptions at finite flow time but do not enlarge the underlying symmetry group of the static lattice regulator beyond the hypercubic subgroup. In this sense DLR realises intrinsic stochastic rotational averaging inside the path integral, rather than an external rotation/projection applied to measured data.
@@
Standard lattice regulators rely on the Symanzik improvement program to systematically remove discretization errors by adding higher-dimension irrelevant operators to the action [16, 17]. On a static hypercubic lattice, the breaking of the Euclidean group $\mathrm{SE}(d)$ down to the hypercubic subgroup $\mathrm{H}(d)$ causes single continuum operators to split into multiple lattice operators, e.g., the Laplacian square $(\partial^{2})^{2}$ splits into rotationally invariant and non-invariant parts. Restoring rotational symmetry to high order therefore requires tuning the coefficients of multiple non-invariant counterterms. In contrast, because DLR preserves exact $\mathrm{SE}(d)$ invariance at finite cutoff, the effective action is constrained to contain only global scalars. This potentially simplifies the improvement program: although the field content is enlarged by the geometry variables, the basis of dangerous irrelevant operators is much smaller. Non-scalar counterterms are forbidden by exact symmetry, meaning one need only tune the coefficients of $\mathrm{E}(d)$-invariant operators to achieve higher-order accuracy.
Moving mesh methods.
From the perspective of numerical analysis, the construction is reminiscent of Arbitrary Lagrangian–Eulerian (ALE) and other moving mesh formulations, in which one works with a fixed reference mesh and a time dependent embedding into physical space, subject to shape regularity constraints [18, 19]. The crucial difference is that in standard ALE methods the mesh motion is prescribed by a deterministic algorithm, often to equidistribute discretization error, and the mesh is not part of the dynamical system. In DLR the embedding variables $x(n)$ are integrated over quantum–mechanically; one may regard the scheme as a “quantum ALE” version of lattice field theory. At a continuum level, one may also view the geometry sector as an “elastic–coordinate” field theory in which one integrates over maps $x:M\to\mathbb{R}^{d}$ on a fixed reference space $M$, with a local energy depending on $\partial x$ and with some nondegeneracy conditions.
These comparisons place DLR as a conservative middle ground between rigid hypercubic lattices and more radical fluctuating–topology approaches: the topology and constructive–QFT structure of Wilson’s lattice theory are retained, while an intrinsic, dynamically fluctuating geometry is used to restore global Euclidean symmetry and, potentially, to reduce lattice artefacts.
2 Geometry of the dynamical lattice regulator